OpenStudy (anonymous):

The quadratic equation x2 + mx + q = 0, where m and q are real numbers, has the complex root a + bi. Which of the following must be equal to m? A.2a B.−2a C.2b D.−2

OpenStudy (anonymous):

it's A or B

OpenStudy (anonymous):

@Aimee98 do you know this

OpenStudy (anonymous):

If α and ß are roots of 1x² + mx + q = 0, then, by Vieta's Formulae, α + ß = -m ........................................... (1) ______________________________ If one root is α = a + bi, then the other root is ß = a - bi. Hence, from (1), ( a + bi ) + ( a - bi ) = -m 2a = -m m = -2a ........................................... Ans.

OpenStudy (anonymous):

B. −2a

OpenStudy (anonymous):

thanks